OpenStudy (bbb911):

__3x - 1___ - ____x - 3____ 2x x Subtract and simplify Will give medal or fanned for the first to help me answer this question :)

OpenStudy (anonymous):

what is that ??

OpenStudy (mathmale):

bbb1996 has presented this problem correctly, except that the fractions are a bit difficult to read. He/she is to combine the two fractions through subtraction and to simplify the resulting equation. bbb: What is the LCD (lowest common denominator) of the two fractions?

OpenStudy (bbb911):

i am thinking its the x

OpenStudy (mathmale):

The x could not be it, because x alone does not include that 2 you see in the other denominator. The LCD MUST include as a factor every distinct quantity in all of the denominators.

OpenStudy (mathmale):

Were you to attempt to add 5/6 and 1/4, you'd need to look at the denominators (6 and 4) and determine the smallest factor that they share. What is it?

OpenStudy (mathmale):

Hints:\[\frac{ 5 }{ 6 }=\frac{ 5 }{ 2*3 },~and~\frac{ 1 }{ 4 }= \frac{ 1 }{ 2*2 }\]do these two denoms. have any common factor?

OpenStudy (mathmale):

(Now waiting for your involvement.)